The sum of the first 6 multiples of 8 , starting with 8 ,is
Found 2 solutions by jim_thompson5910, vleith:
Answer by jim_thompson5910(35256)

8+16+24+32+40+48=168

Answer by vleith(2983)

1*8 + 2*8 + 3*8 + 4*8 + 5*8 + 6*8 = (1+2+3+4+5+6) * 8 = 21*8 = 168